how do you find the domain in interval notation of 3x^2-7x+5=11
Answer by shree840(260)

I believe you get the answer of x by solving the quadratic eequation
3X^2-7x+5=11
3x^2-7x-6=0
3x^2-9x+2x-6=0 remember you need to multiply 3 with -6 and find the factors
which are -9 +2
3x(x-3)+2(x-3)
(3x+2)(x-3)
x=-2/3 or 3 ANS